OpenStudy (slayer8001):

Find the radical equivalent of 27^2/3 and compute it.

OpenStudy (anonymous):

hang on

OpenStudy (anonymous):

3√27^2

OpenStudy (slayer8001):

@Vampirelove28 you literally just got that off of algebra.com and that's not even the same equation that ones fraction is negative

OpenStudy (jango_in_dtown):

radical means expressions having square root or cube root or... see27^2/3=cube root of 27^2, which is the radical equivalent

OpenStudy (slayer8001):

h(x) = 4^x + x + 6 h(x) = 4^x + 1 -2 h(x) = 4^x + x - 2 h(x) = 4^x - 4

OpenStudy (jango_in_dtown):

now the second part, we know that 27=3^3 so we have cube root of 27^2=((3^3)^2)^(1/3)=3^(6/3)=3^2=9

OpenStudy (jango_in_dtown):

@slayer8001